OpenStudy (anonymous):

Solve the equation 2(h – 8) – h = h – 16 A)h = 8 B)h = –8 C)h = infinitely many solutions D)h = no solution

OpenStudy (anonymous):

c

OpenStudy (anonymous):

2h - 16 - h = h - 16 h - 16 = h - 16 C